OpenStudy (anonymous):

how can we learn into function onto function etc, these are too difficult to learn.... any one who help me?

OpenStudy (anonymous):

psot a specific example

OpenStudy (anonymous):

actually i have no basic concepts abut..it then how can i give examp;e

OpenStudy (anonymous):

A one-to one mapping, also known as an injection, is a mapping from set A to set B such that for all b in B there exists at most one a in A such that b = f(a). An onto mapping, also known as a surjection, is a mapping of a set A to set B such that for every b in B there exists at least one a in A such that b = f(a). A one-to-one onto mapping, a.k.a. one-to-one correspondence, a.k.a. bijection, is a mapping from A to B such that for every b in B there exists exactly one a in A such that b = f(a), and for every a in A there exists exactly one b in B such that f(a) = b.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

To test the graph of a function for one-to-one, draw horizontal oines. If all horizontal lines cross the graph at no more than one point, then the function is one-to-one. But if even one horizontal line crosses the graph at two or more points, then the function is not one-to-one. To test for an onto function, imagine vertical lines drawn through each point in the domai, All those lines must intersect the graph of the functiom at exactly one point. If any vertical line through a point in the domain does not intersect the graph, then the function is not an onto function.